OpenStudy (anonymous):

Suppose cos A = 12/13 with 0º≤A≤90º. Suppose also that sin B = 8/17 with 90º≤B≤180º. Find tan (A + B). I think I found the correct tan A and tan B Tan A= 5/12 and tan B=-8/15 (correct me if I'm wrong) How would I find tan (A+B)?

OpenStudy (anonymous):

tan (A+B) = (tan A + tan B)/ (1-tanAtanB) = (5/12 - 8/15)/(1 - (5/12)(8/15)) The rest is arithmetic.

OpenStudy (anonymous):

You are correct for the values of tan A and tan B. so we just substituted those values into the formula.

OpenStudy (anonymous):

so how would i simplify this? i was never good with fractions.